archive-es.com » ES » O » OVH.ES

Total: 788

Choose link from "Titles, links and description words view":

Or switch to "Titles and links view".
  • OVH Tasks  
    Network Task Type Maintenance Category the whole network Status CLOSED Percent Complete Details The link Paris Madrid will be interrupted the 16th of May at 00 00 PM EST for the fiber rerouting The estimated disconnection time is 4h Date

    Original URL path: http://status.ovh.es/?do=details&id=37 (2015-10-15)
    Open archived version from archive


  • OVH Tasks  
    Category the whole network Status CLOSED Percent Complete Details A software maintenance is successively estimated on our ports T SYSTEMS Paris and Frankfurt on the 3rd of May 2010 between 3 00 ET 5 30 PM EST The traffic will

    Original URL path: http://status.ovh.es/?do=details&id=38 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    change Date Thursday 06 May 2010 15 25PM Reason for closing Done Comments 23 Comment by OVH Tuesday 04 May 2010 10 45AM We start with server 1 A stopping for a few minutes Comment by OVH Tuesday 04 May 2010 10 46AM Chassis change completed data import Comment by OVH Tuesday 04 May 2010 10 48AM Private sqls of server1 restart verification of good functioning Comment by OVH Tuesday 04 May 2010 10 50AM The intervention is ok for server1 everything has been restarted and is functional We pass to server2 a stopping in a few minutes Comment by OVH Tuesday 04 May 2010 10 50AM server2 is switched reboot is ongoing Comment by OVH Tuesday 04 May 2010 10 53AM Around server3 a stopping is ongoing Comment by OVH Tuesday 04 May 2010 10 53AM data import completed reboot is ongoing Comment by OVH Tuesday 04 May 2010 10 53AM Private sqls are restarted everything is ok on server3 Comment by OVH Tuesday 04 May 2010 10 55AM The rest will be on next night Comment by OVH Tuesday 04 May 2010 10 55AM We start this night with server4 Stopping of private sqls in a few minutes Comment by OVH Tuesday 04 May 2010 10 56AM Data transfer completed for server4 We reboot the private sqls Comment by OVH Tuesday 04 May 2010 10 56AM Around server 15 we turn it off and we transfer data to the new chassis Comment by OVH Tuesday 04 May 2010 10 56AM Transfer completed we reboot server 15 Comment by OVH Tuesday 04 May 2010 10 57AM We turn off server5 Comment by OVH Tuesday 04 May 2010 10 57AM transfer of server5 completed we reboot it Comment by OVH Tuesday 04 May 2010 10 57AM extinction of server6 Comment

    Original URL path: http://status.ovh.es/?do=details&id=39 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Percent Complete Details GSW proceeds to a periodical electrical maintenance on the A way on this Tuesday 04 05 2010 between 9h and 12h30 Each one of our bays are supplied with 2 ways A and B via an STS

    Original URL path: http://status.ovh.es/?do=details&id=40 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Category all dedicated servers Status CLOSED Percent Complete Details The server that manages among others MRTG graphs of networks 94 23 40 xx to 94 23 47 xx displays some signs of weakness We are currently diagnosing the problem The

    Original URL path: http://status.ovh.es/?do=details&id=42 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    kernels If you have an existing Ubuntu installation we strongly recommend you not to upgrade to Ubuntu 10 04 LTS yet Both Desktop and Server versions are affected We are working on a way to circumvent this issue Date Wednesday 08 September 2010 13 36PM Reason for closing Done Comments 3 Comment by OVH Tuesday 08 June 2010 19 50PM The Ubuntu Server 10 04 LTS Lucid Lynx distribution has been made available for installations via Manager For the time being the Ubuntu kernel is used If you choose to upgrade your existing Ubuntu installation please keep in mind that it won t boot on any kernel except the original Ubuntu kernel a manual upgrade is still not recommended We will update this task with new information as available Comment by OVH Tuesday 08 June 2010 19 50PM The according kernel configuration options have been identified the new configurations are now being tested for backwards compatibility Comment by OVH Wednesday 23 June 2010 17 45PM New compatible kernels are now available for dedicated servers GRS kernels ftp ftp ovh net made in ovh bzImage 2 6 33 5 2 STD kernels ftp ftp ovh net made in ovh bzImage 2

    Original URL path: http://status.ovh.es/?do=details&id=43 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Percent Complete Details We have just finished the interconnection with a new supplier We will switch the outgoing traffic on this link No disruption is expected No perturbation is estimated New calls will pass through and ongoing calls will continue

    Original URL path: http://status.ovh.es/?do=details&id=45 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Percent Complete Details GSW proceeds to a periodical electrical maintenance on the B way on this Thursday 06 05 2010 between 9h and 12h30 Each one of our bays are supplied with 2 ways A and B via an STS

    Original URL path: http://status.ovh.es/?do=details&id=41 (2015-10-15)
    Open archived version from archive